OpenStudy (anonymous):

Please rearrange to make y the subject. x=2/3y^2+4

OpenStudy (anonymous):

\[y=\sqrt{2-4x/3x}\]

OpenStudy (anonymous):

can u write it clearly by equation tag it would be much easier if u had not understand the whole solution

OpenStudy (anonymous):

\[2/3y ^{2}+4=x\]

OpenStudy (anonymous):

maybe mathlearners 3 and 2 should be swapped around?

OpenStudy (anonymous):

i think my answer is wrong. i think this is the right one.... \[\sqrt{2/-4+3x}\]

OpenStudy (anonymous):

\[y=\sqrt{3/2x-4}\] is the right answer!

OpenStudy (anonymous):

sorry:D

OpenStudy (anonymous):

its all good

OpenStudy (anonymous):

how did it become \[√y=3/2x−4\]